# From Frames to Clips: Efficient Key Clip Selection for Long-Form Video Understanding

### v1

@@ 5 Experiments
Table 3: Ablation study on Video-MME.
Method Short Medium Long Overall
Uniform 67.3 55.0 48.9 57.1
Watershed 70.0 58.3 50.7 59.7
+ Key Clip (Fixed) 73.0 60.0 53.7 62.2
+ Key Clip (Adaptive) 73.8 62.2 54.1 63.4
Table 4: Comparison of temporal modeling on Video-MME. Baselines use 64 isolated frames, while F2C uses 16 clips of 4 frames.
Method # Frames Short Medium Long Overall
Uniform 64 71.3 60.1 51.4 61.0
Top-$k$ 64 73.6 59.9 52.3 61.9
BOLT 64 73.4 61.0 50.6 61.7
Watershed 64 72.3 59.8 53.4 61.9
[…]
5.1 Experimental Settings
Benchmarks. We evaluate our method on three widely used long-form video benchmarks: Video-MME, LongVideoBench, and MLVU, which together cover diverse domains and evaluation protocols. Video-MME (Fu et al., 2024) is a large-scale multimodal evaluation suite designed to assess the capabilities of Video LLMs across video understanding, reasoning, and knowledge-grounded tasks. It includes both short and long-form videos, enabling a comprehensive evaluation of context management strategies. LongVideoBench (Wu et al., 2024b) specifically targets long-form video understanding with a diverse set of tasks such as VQA, summarization, and temporal grounding. It emphasizes scenarios where videos span tens of minutes to hours, providing a challenging testbed for evaluating temporal reasoning and efficient context construction. MLVU (Multi-level Long Video Understanding) (Zhou et al., 2024) is a benchmark focused on hierarchical understanding of long videos. It evaluates models on multiple levels of reasoning, including frame-level recognition, clip-level temporal understanding, and video-level holistic comprehension. This makes MLVU particularly suitable for testing whether methods like ours preserve both local and global context.
Comparison Methods. We compare our proposed method with several representative selection strategies which are applicable to long-form videos. Uniform sampling selects frames at equal temporal intervals across the video. Top-$k$ sampling selects the frames with the highest similarity scores to the given text query. The training-free method BOLT (Liu et al., 2025) selects frames using Inverse Transform Sampling based on frame importance scores. Q-frame (Zhang et al., 2025) selects three levels of frames in different resolutions. In addition, we include watershed selection, which selects frames according to temporal boundaries detected by watershed segmentation.
Implementation Details. We use Qwen2.5-VL-7B as the backbone VLM due to its support for any input resolution. All videos are loaded at 1 FPS for selection. For all baseline methods, we keep the original video resolution without resizing to ensure a fair comparison. Evaluations are conducted using the lmm_evals library (Zhang et al., 2024b) on a computing cluster equipped with NVIDIA A100 GPUs. For similarity-based selection (Top-$k$), we use SigLIP2 (Tschannen et al., 2025) as the vision-language model to compute similarity scores. More details are provided in Appendix D.
5.2 Results on Long-form Video Benchmarks
Table 2 summarizes the results on Video-MME, LongVideoBench, and MLVU under different frame budgets. We compare F2C with uniform sampling, Top-$k$, watershed, Q-Frame, and BOLT.
Across all the three benchmarks, F2C consistently surpasses baselines. Compared with uniform sampling, it yields substantial improvements, with the largest gains on MLVU, followed by Video-MME, and smaller but steady gains on LongVideoBench. The benefit is most pronounced under small frame budgets (e.g., $K=8$), where uniform sampling often misses critical content, while F2C remains robust by preserving temporal continuity. As $K$ increases, the relative gap narrows, but F2C still retains a stable advantage.
Figure 5: Impact of different initial keyframe selectors on Video-MME with $K=16$.
@@ 6.1 Ablation Study
6.2 Impact of Temporal Information
To examine whether temporal continuity provides greater benefit than simply sampling more frames, we conduct an experiment with a total of 64 frames. For baselines, 64 isolated frames are selected directly, while in F2C the same budget is allocated to 16 clips with up to 4 frames each. To ensure fairness, all frames are downsampled with $s=2$, so the token count remains comparable across methods. Results are reported in Table 4.
F2C consistently outperforms the baselines, with the largest gain on long videos, where it reaches $54.1$ compared to 50.650.653.453.4 for separate-frame approaches. This demonstrates that grouping frames into clips yields richer temporal cues than simply increasing frame count. Moreover, since F2C introduces resolution as an additional factor in context construction, the results also indicate that resolution-aware design can further enhance existing frame selection methods.
@@ 6.4 Computational Analysis
We analyze the computational efficiency of F2C by comparing visual token counts on Video-MME (w/o subtitles) under budgets of 8, 16, and 32 full-resolution frames (Table 5). Despite adding temporal context, F2C achieves similar or fewer tokens than frame-based selectors because overlapping frames are encoded only once. The reduction becomes increasingly significant as the budget grows, showing that F2C not only enriches temporal continuity but also improves efficiency. This balance of accuracy and computation makes F2C well suited for long-form video understanding.
@@ 6.5 Visualization
Figure 7 provides a qualitative comparison of different frame selection strategies on a long-form video. Uniform sampling and BOLT fail to capture the crucial frames corresponding to the correct answer, resulting in incorrect predictions. Top-K manages to select one relevant frame but lacks sufficient coverage of the question, leading to incomplete context. In contrast, our method successfully selects the key frames related to the correct option while also maintaining high diversity across the video, enabling more comprehensive temporal reasoning and accurate prediction. Besides, we provide a visualization on the distribution of the selections in Appendix E.
Figure 7: Visualization of the selected frames for each method.
@@ 7 Conclusion
We revisited the frame selection for long-form video understanding and proposed Frames-to-Clips, a training-free framework that replaces isolated keyframes with temporally coherent clips and introduces adaptive trade-off between resolution and clip length under a fixed token budget. Extensive experiments on Video-MME, LongVideoBench, and MLVU show that F2C consistently outperforms state-of-the-art training-free baselines. Our analyses further demonstrate that temporal continuity, diversity in anchor selection, and balanced resolution scaling are all crucial for effective context management. By improving selection without training, F2C offers a simple and scalable solution for long-form video understanding and points toward future directions for VLMs.
Appendix
@@ Appendix A Overview
  • Visualization of Selection Distributions (Section E): illustrates qualitative comparisons of frame selection strategies.
@@ Appendix C Limitation
Although F2C effectively improves context management by providing temporally coherent and resolution-adaptive inputs, its performance upper bound is still constrained by the capability of the downstream VLM. Even when the correct frames or clips are selected, the final answer depends on the reasoning and comprehension ability of the VLM itself. As a result, limitations in temporal reasoning, spatial understanding, or multimodal alignment within the backbone model remain bottlenecks. Our method is thus complementary to future advances in Video LLM architectures, and its benefits may further amplify as more powerful backbone models become available.
@@ D.3 Hyperparameter of F2C
D.4 Re-implementation of BOLT and Q-Frame
For a fair comparison, we closely followed the official settings reported in the original papers. For BOLT (Liu et al., 2025), we adopted the same hyperparameters as ours and set $\alpha=2.5$ as recommended in their work. For Q-Frame (Zhang et al., 2025), the original implementation evaluates three configurations of $(\text{num}_{\text{high}},\text{num}_{\text{medium}},\text{num}_{\text{low}})=(4,8,32)$, which corresponds to our 8-frame setting. To extend Q-Frame under larger budgets, we applied proportional scaling and used $(8,16,64)$ for the 16-frame setting and $(16,32,128)$ for the 32-frame setting. These adjustments ensure that both baselines are re-implemented consistently and evaluated within our experimental framework, facilitating fair comparison and reproducibility.
Appendix E Visualization on Distribution of Selection
@@
Uniform sampling spreads frames evenly but ignores semantic signals, often missing important peaks. BOLT tends to favor high-similarity regions but suffers from redundancy. Top-$k$ focuses too narrowly, selecting clustered frames around peaks while overlooking other segments. In contrast, F2C selects temporally coherent clips that cover diverse regions with high relevance, striking a better balance between coverage and precision.
@@ F.1 Impact of the Version of CLIP model
LongCLIP (Zhang et al., 2024a) 248 61.9 63.1 65.1
CLIP (Radford et al., 2021) 77 61.9 65.1 65.7
SigLIP (Zhai et al., 2023) 64 62.5 64.8 65.4
SigLIP2 (Tschannen et al., 2025) 64 61.4 63.4 65.6
Figure 8: Visualization of the selected frames

### v2

@@ 5 Experiments
Figure 5: Impact of different initial keyframe selectors on Video-MME with $K=16$ and extending them to key clips by F2C.
@@
5.1 Experimental Settings
Benchmarks. We evaluate our method on three widely used long-form video benchmarks: Video-MME, LongVideoBench, and MLVU, which together cover diverse domains and evaluation protocols. Video-MME [fu2024video] is a large-scale multimodal evaluation suite designed to assess the capabilities of Video LLMs across video understanding, reasoning, and knowledge-grounded tasks. It includes both short and long-form videos, enabling a comprehensive evaluation of context management strategies. LongVideoBench [wu2024longvideobench] specifically targets long-form video understanding with a diverse set of tasks such as VQA, summarization, and temporal grounding. It emphasizes scenarios where videos span tens of minutes to hours, providing a challenging testbed for evaluating temporal reasoning and efficient context construction. MLVU (Multi-level Long Video Understanding) [zhou2024mlvu] is a benchmark focused on hierarchical understanding of long videos. It evaluates models on multiple levels of reasoning, including frame-level recognition, clip-level temporal understanding, and video-level holistic comprehension. This makes MLVU particularly suitable for testing whether methods like ours preserve both local and global context.
Comparison Methods. We compare our proposed method with several representative selection strategies that are applicable to long-form videos. Uniform sampling selects frames at equal temporal intervals across the video. Top-$k$ sampling selects the frames with the highest similarity scores to the given text query. The training-free method BOLT [BOLT] selects frames using Inverse Transform Sampling based on frame importance scores, AKS [aks] adaptively selects keyframes by recursively splitting video segments based on frame–text relevance, and Q-frame [zhang2025qframequeryawareframeselection] selects three levels of frames in different resolutions. In addition, we include watershed selection, which selects frames according to temporal boundaries detected by watershed segmentation.
Implementation Details. We use Qwen2.5-VL-7B as the backbone VLM due to its support for any input resolution. All videos are loaded at 1 FPS for selection. For all baseline methods, we keep the original video resolution without resizing to ensure a fair comparison. Evaluations are conducted using the lmm_evals library [zhang2024lmmsevalrealitycheckevaluation] on a computing cluster equipped with NVIDIA A100 GPUs. For similarity-based selection (Top-$k$), we use SigLIP2 [tschannen2025siglip2multilingualvisionlanguage] as the vision-language model to compute similarity scores. More details are provided in the supplementary D.
5.2 Results on Long-form Video Benchmarks
Table 2 summarizes the results on Video-MME, LongVideoBench, and MLVU under different frame budgets. We compare F2C with uniform sampling, Top-$k$, watershed, BOLT, AKS, and Q-Frame.
Across all benchmarks, F2C consistently surpasses baselines. Compared with uniform sampling, it yields substantial improvements, with the largest gains on MLVU, followed by Video-MME, and smaller but steady gains on LongVideoBench. The benefit is most pronounced under small frame budgets (e.g., $K=8$), where uniform sampling often misses critical content, while F2C remains robust by preserving temporal continuity. As $K$ increases, the relative gap narrows, but F2C still retains a stable advantage.
@@ 6 Discussion
Figure 7: Visualization of the selected frames for each method.
@@ 6.1 Ablation Study
6.2 Impact of Temporal Continuity
To examine whether temporal continuity provides greater benefit than simply sampling more frames, we conduct an experiment with a total of 64 frames. For baselines, 64 isolated frames are selected directly, while in F2C, the same budget is allocated to 16 clips with up to 4 frames each. To ensure fairness, all frames are downsampled with $s=2$, so the token count remains comparable across methods. Results are reported in Table 4.
F2C consistently outperforms the baselines, with the largest gain on long videos, where it reaches $54.1$ compared to 50.650.653.453.4 for separate-frame approaches. This demonstrates that grouping frames into clips yields richer temporal cues than simply increasing frame count. Moreover, by introducing resolution as an additional factor in context management for long-form video understanding, such a resolution-aware design can further enhance existing frame selection methods.
@@ 6.4 Computational Analysis
We analyze the computational efficiency of F2C by comparing visual token counts on Video-MME (w/o subtitles) under budgets of 8, 16, and 32 full-resolution frames (Table 5). Despite adding temporal context, F2C achieves similar or fewer tokens than frame-based selectors because overlapping frames are encoded only once. The reduction becomes increasingly significant as the budget grows, showing that F2C not only enriches temporal continuity but also improves efficiency.
It is also worth noting that the additional cost introduced by the vision-language encoder is minor relative to the overall VLM inference cost. First, image features can be pre-computed and cached, making the per-query overhead effectively zero during inference. Second, even when computed on the fly, the encoding is parallel across frames and can be executed in a single batch, whereas VLM decoding remains strictly sequential over tokens. As a result, the dominant computational bottleneck remains the LLM forward pass rather than the visual encoder. This balance of accuracy and computation further underscores the suitability of F2C for long-form video understanding.
@@ 6.5 Visualization
Figure 7 provides a qualitative comparison of different frame selection strategies on a long-form video. Uniform sampling and BOLT fail to capture the crucial frames corresponding to the correct answer, resulting in incorrect predictions. Top-K manages to select one relevant frame but lacks sufficient coverage of the question, leading to incomplete context. In contrast, our method successfully selects the key frames related to the correct option while also maintaining high diversity across the video, enabling more comprehensive temporal reasoning and accurate prediction. Besides, we provide a visualization on the distribution of the selections in the supplementary E.
@@ 7 Conclusion
We revisited the frame selection for long-form video understanding and proposed Frames-to-Clips, a training-free framework that replaces isolated keyframes with temporally coherent clips and introduces an adaptive trade-off between resolution and clip length under a fixed token budget. Extensive experiments on Video-MME, LongVideoBench, and MLVU show that F2C consistently outperforms state-of-the-art training-free baselines. Our analyses further demonstrate that temporal continuity, diversity in anchor selection, and balanced resolution scaling are all crucial for effective context management. By improving selection without training, F2C offers a simple and scalable solution for long-form video understanding and points toward future directions for VLMs.
[…]
Supplementary Material
@@ Appendix A Overview
  • Visualization (Section E): illustrates qualitative comparisons of frame selection strategies and cases analysis on F2C.
@@ Appendix C Limitation
Although F2C effectively improves context management by providing temporally coherent and resolution-adaptive inputs, its performance upper bound is still constrained by the capability of the downstream VLM. Even when the correct frames or clips are selected, the final answer depends on the reasoning and comprehension ability of the VLM itself. As a result, limitations in temporal reasoning, spatial understanding, or multimodal alignment within the backbone model remain bottlenecks. Our method is thus complementary to future advances in Video LLM architectures, and its benefits may further amplify as more powerful backbone models become available. A further discussion and visualization are provided in Section E.2.
@@ D.3 Hyperparameter of F2C
D.4 Re-implementation of BOLT, Q-Frame, and AKS
For a fair comparison, we closely followed the official settings reported in the original papers. For BOLT [BOLT], we adopted the same hyperparameters as ours and set $\alpha=2.5$ as recommended in their work. For Q-Frame [zhang2025qframequeryawareframeselection], the original implementation evaluates three configurations of $(\text{num}_{\text{high}},\text{num}_{\text{medium}},\text{num}_{\text{low}})=(4,8,32)$, which corresponds to our 8-frame setting. To extend Q-Frame under larger budgets, we applied proportional scaling and used $(8,16,64)$ for the 16-frame setting and $(16,32,128)$ for the 32-frame setting. For AKS [aks], we integrate the official implementation (t1=$0.1$, t2=100-100, and all_depth=5) into our framework with the same SigLIP2 feature extractor as other methods for a fair comparison on frame selection. These adjustments ensure that both baselines are re-implemented consistently and evaluated within our experimental framework, facilitating fair comparison and reproducibility.
Appendix E Visualization
E.1 Distribution of Selection
@@
Uniform sampling spreads frames evenly but ignores semantic signals, often missing important peaks. BOLT tends to favor high-similarity regions but suffers from redundancy. Top-$k$ and AKS focus too narrowly, selecting clustered frames around peaks while overlooking other segments. In contrast, F2C selects temporally coherent clips that cover diverse regions with high relevance, striking a better balance between coverage and precision.
@@
E.2 Selected Frames of F2C
Figure 9 illustrates that F2C effectively captures key frames with coherent temporal progression, which is particularly beneficial for action-related VQA. Figure 10 demonstrates F2C’s ability to adaptively balance clip length and spatial resolution based on content, while Figure 11 shows that the method also performs well on questions requiring high-level contextual understanding.
We additionally present representative failure cases. In Figure 12, F2C successfully identifies relevant frames, but the downstream VLM fails to reason correctly, leading to an incorrect answer. In Figure 13, for more complex reasoning-oriented questions, F2C does not fully capture the critical event, highlighting potential limitations when the key evidence is extremely sparse or subtle.
Figure 8: Visualization of frame selection distributions for different methods. The similarity curve denotes the frame-level relevance to the query. Uniform sampling distributes frames evenly but ignores semantic peaks; BOLT concentrates on high-similarity regions yet exhibits redundancy; Top-$k$ and AKS select excessively clustered frames with limited coverage. In contrast, F2C selects temporally coherent clips that cover diverse, high-relevance regions while maintaining temporal continuity, leading to more informative and efficient context construction.
@@ F.1 Impact of the Version of CLIP model
LongCLIP [zhang2024longclip] 248 61.9 63.1 65.1
CLIP [radford2021learning] 77 61.9 65.1 65.7
SigLIP [zhai2023sigmoid] 64 62.5 64.8 65.4
SigLIP2 [tschannen2025siglip2multilingualvisionlanguage] 64 61.4 63.4 65.6
[…]
Figure 12: Failure case in Video-MME. Although F2C successfully selects frames highly relevant to the query, the downstream VLM fails to perform the required reasoning, resulting in an incorrect prediction. This example illustrates that overall performance remains bounded by the reasoning capability of the underlying VLM, even when key visual evidence is correctly captured.
Figure 13: Failure Case: Key Evidence Not Fully Captured. In this example, the critical event appears only briefly within the long video, and F2C does not select the decisive frames required to answer the question correctly. Although many relevant clips are included, the key moment is too sparse and subtle, highlighting a limitation of clip-based selection when essential visual cues are extremely short-lived.